A sample of mass 5g of an unknown gas produced a pressure of 19.4kPa in a 10L flask at 25oC. Calculate the molar mass of the gas.
1
Expert's answer
2020-09-21T06:30:29-0400

Solution.

"p \times V = \frac{m}{M} \times R \times T"

"M = \frac{m \times R \times T}{p \times V}"

"M = \frac{0.005 \times 8.31 \times 298}{19400 \times 0.01} = 0.06382 \ kg/mol = 63.82 \ g/mol"

M = 63.82 g/mol

Answer:

M = 63.82 g/mol